# AdvRepair: Provable Repair of Adversarial Attack

### v1

@@ 4.4 Scalability
4.5 Monotonicity w.r.t. the number of buggy behaviors
We systematically evaluate the performance of the repair results concerning the number of buggy behaviors in the local robustness repair task. We consider five distinct configurations: 50, 100, 200, 500 and 1000. The comparative analysis encompasses the performance of various baselines alongside our approach across these configurations. AdvRepair not only showcases effectiveness in limiting the number of buggy behaviors but also excels when faced with a growing number of buggy behaviors. In contrast, APRNN encounters challenges in successfully repairing the FNN_small model when the number of buggy behaviors exceeds 100, and REASSURE experiences memory overflow issues when confronted with 1000 buggy behaviors.
Answer to RQ4: Even in scenarios involving a large number of buggy behaviors, AdvRepair demonstrates efficient and effective removal of buggy behaviors while preserving the original performance of the models. Comparative analysis indicates that AdvRepair leverages the available buggy points to achieve superior repair results compared to alternative approaches.
5 Related work
@@
Provable DNN repair. Our approach falls within the category of provable DNN repair methods. The methods most closely related to ours are ART [39] and REASSURE [22]. In comparison to REASSURE, our approach fundamentally differs in its repair objectives. REASSURE focuses on fixing activation patterns of neural networks, while we directly guarantee the properties, whose input constraints may encompass a large number of activation patterns. In this way, our method exhibits superior generalization compared to REASSURE. Although ART also utilizes training in its repair process, it directly tunes parameters in the original neural network. In contrast, our approach involves training patch modules specifically for different properties, enhancing the specialization of our repairs and, consequently, their effectiveness. Other methods such as PRDNN [60] and APRNN [63] formulate the repair problem in linear programming. However, their effectiveness diminishes in properties with high-dimensional polytopes, such as the robustness of image classifications.
@@
Heuristic DNN repair. Utilizing heuristic algorithms such as particle swarm optimization and differential evolution, CARE [61] and Arachne [59] aim to localize the neurons responsible for the fault. VeRe [41] focuses on providing formal verification guidance for fault localization and defining target intervals for repair synthesis. DeepRepair [82] and few-shot guided mix [51], for instance, expand the set of negative samples to generate additional training data. Tian et al. [64], on the other hand, augment the existing data by introducing real-world environmental effects, like fog, to the samples. To the best of our knowledge, DL2 [21] is the first paper to combine logical constraints and loss function, but it does not adopt convex approximation to give the loss function a certified guarantee.
@@
DNN verification. In 2010, [50] proposed the first method for DNN verification. Over the past decade, various approaches of DNN verification has been proposed including techniques such as constraint solving [9, 25, 31, 38, 47, 33, 18, 28], abstract interpretation [24, 57, 80, 56, 36, 55], linear relaxation [30, 49, 79, 74, 7], global optimisation [52, 17, 53], CEGAR [2, 19, 48], reduction to two-player games [75, 77], and star-set abstraction [67, 66]. These method offer provable estimations of DNN robustness. Moreover, statistical approaches, presented in [72, 73, 5, 76, 6, 11, 27, 44, 12, 37], prove to be more efficient and scalable, particularly suited for intricate DNN structures, allowing for the establishment of quantifiable robustness at a specified confidence level. Certified training [4, 45, 46] uses convex approximation in training the loss function, but it is only used to enhance the robustness of DNN, without making it accessible to repair properties in the training procedure.
6 Conclusion
@@
We introduce AdvRepair, a novel approach for addressing adversarial attacks using limited data. Our method provides provable and targeted repairs within the robustness neighborhood, enabling the generalization of this defense to other inputs. In terms of efficiency, scalability, and generalization, our approach surpasses existing methods. For future work, we plan to integrate BaB-based verification methods to enhance the performance of our approach, improving both precision and offering an alternative refinement approach. Additionally, our patch-based repair framework has the potential to evolve into a black-box repair method.
Appendix 0.A Proof of Theorem 3.4
Theorem 0.A.1
Let $\varphi=(N,B(x_{i},r))$ be a local robustness property. If $\mathcal{L}(\varphi)=0$ on $B(x_{i},r)$, i.e.,
$\mathcal{L}^{*}(\varphi):=\max(\max(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}+r\cdot\bm{1})+\min(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}-r\cdot\bm{1})+\beta_{\ell},0),$
@@
where $\bm{0}$ and $\bm{1}$ are the vector in $\mathbb{R}^{n_{0}}$ with all the entries 00 and $1$, respectively, then the property $\varphi$ holds.
Proof
@@
$\max_{x\in B(x_{i},r)}\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$
$=$ $\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}>0\}}(\alpha_{\ell})_{i}\max_{x\in B(x_{i},r)}x+\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}<0\}}(\alpha_{\ell})_{i}\max_{x\in B(x_{i},r)}x+\beta_{\ell}$
$=$ $\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}>0\}}(\alpha_{\ell})_{i}\cdot(x_{i}+r)+\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}<0\}}(\alpha_{\ell})_{i}\cdot(x_{i}-r)+\beta_{\ell}$
@@
$\iff$ $\sum_{\ell\neq\ell_{0}}\max(\max(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x+r\cdot\bm{1})+\min(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x-r\cdot\bm{1})+\beta_{\ell},0)=0$
@@
$\forall x\in B(x_{i},r),C_{N}(x)_{\ell}-C_{N}(x)_{\ell_{0}}\leq\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}\leq 0,$ (2)
@@
Appendix 0.B Impact of Patch Module Size and Quantity on Efficiency
In this experiment, we assess the efficacy of size and quantity of patch modules in the context of repairing local robustness properties. We conduct a comparative analysis of AdvRepair’s performance across various sizes and quantities, presenting the experimental outcomes in Table 5. Two distinct patch module sizes, namely small and large, are considered. The small patch module consists of three hidden layers with 64 neurons in the first two layers and 16 neurons in the last layer, while the large patch module comprises three hidden layers with 256 neurons in the first two layers and 64 neurons in the last layer. Additionally, we explore two different numbers of patch modules: property-guided patch modules and untargeted label patch modules.
Table 5: The efficacy of patch module scale and quantity
Model $r$ $n$ RSR/RGR/DSR/DGSR/%
PS PL LS LL PS PL LS LL PS PL LS LL PS PL LS LL
FNN_ small 0.05 50 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 90.0 90.0 89.9 89.8
100 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.4 96.4 96.1 96.0
200 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.6
500 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.6
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.6
0.1 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.3 96.3 89.5 89.5
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 95.0 96.3 94.7 94.8
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 95.2 95.1 94.9 94.8
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 95.3 95.2 96.3 96.3
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.5 96.5 96.3 96.3
0.3 50 94.0 94.0 94.0 94.0 94.8 94.8 94.8 94.8 92.0 92.0 92.0 92.0 86.4 79.9 44.6 44.5
100 99.0 99.0 99.0 99.0 99.0 99.0 99.0 99.0 100.0 99.0 99.0 100.0 96.2 96.2 85.8 86.1
200 99.5 99.5 99.5 99.5 99.5 99.5 99.5 99.5 100.0 99.0 97.5 98.0 90.8 91.2 85.9 85.9
500 99.4 99.4 99.8 99.8 99.5 99.5 99.8 99.8 99.8 99.6 100.0 100.0 92.0 92.1 89.5 89.7
1000 99.6 99.6 99.9 99.9 99.7 99.7 100.0 100.0 99.7 99.6 100.0 100.0 96.6 96.5 94.2 94.2
FNN_ big 0.05 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.0 97.0 96.8 96.8
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.8 96.8
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.8 96.8
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 97.2 97.2
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 97.2 97.2
0.1 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.8 96.8
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.2 96.8 96.9
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.9 96.8
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 96.8 96.8
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 97.2 97.0
0.3 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.0 96.1 89.1 89.1
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.7 96.8 89.1 89.1
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 89.3 89.3
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 95.9 95.9
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.9 99.9 100.0 100.0 97.2 97.2 97.1 97.1
CNN 0.05 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.7
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
0.1 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.2 98.2 98.0 98.0
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.2 98.2 98.1 98.0
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.5 99.5 99.5 98.2 98.2 98.0 98.1
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.8 99.8 99.8 99.8 98.2 98.2 98.0 98.1
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.9 99.8 99.8 99.8 98.3 98.3 98.0 98.0
0.3 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.2 98.2 98.3 98.3
The performance of AdvRepair is evaluated across different combinations of patch module scales and quantities, including small property-guided patch module (PS), large property-guided patch modules (PL), small untargeted label repairing patch modules (LS), and large untargeted label repairing patch modules (LL). Notably, the experimental findings indicate that both sizes of the patch modules exhibit commendable performance, with minimal impact observed on the repair performance due to the network size. Concerning the quantity of patch modules, the property-guided patch modules outperform the untargeted label repairing patch modules in defending against adversarial attacks within $D_{t}$. We hypothesize that this superiority is attributed to the specialization inherent in property-guided repair.
Appendix 0.C Experimental Details

### v2

@@ IV-B Repair performance
IV-C Generalization
In the experiment, we assess the repaired network’s generalization and resistance to new adversarial attacks (Table III). PatchPro leads in $\mathrm{RGR}$, demonstrating superior performance compared to other baselines. PRDNN, APRNN, and TRADES also show some degree of generalization, particularly on CNNs. For $\mathrm{DSR}$, PatchPro outperforms competitors, with TRADES showing moderate defense against small-radius attacks on FNN_big and CNN. Evaluating $\mathrm{DGSR}$, PatchPro again tops the list, while CARE, PRDNN, APRNN, and TRADES show limited defense against FNN_big and CNN. $\mathrm{DGSR}$ is key for gauging generalization to global inputs . Overall, these findings underscore PatchPro’ breakthrough in repairing adversarial attacks.
Answer to RQ1 and RQ2: PatchPro consistently outperforms the baselines in both local robustness repair and safety property violation correction tasks. It achieves 100% repair success rate coupled with 0% drawdown and acceptable efficiency. Additionally, it demonstrates significant generalization against unforeseen adversarial attacks.
IV-D Scalability Evaluation
We examine PatchPro’ scalability on VGG19 and ResNet18 for CIFAR-10 local robustness repair via feature-space repair. Results in Table IV show PatchPro, PRDNN, and APRNN achieve 100% $\mathrm{RSR}$, while TRADES follows closely. PatchPro leads in $\mathrm{RGR}$ and $\mathrm{DD}$ with TRADES following suit. PatchPro also significantly outperforms others in $\mathrm{DSR}$ and matches accuracy levels in $\mathrm{DGSR}$, surpassing all tools. On larger datasets like Tiny ImageNet, PatchPro maintains superiority in repairing WRN101-2 and ResNet152, as seen in Table V. PatchPro’ running time is reasonable given its superior repair and generalization abilities. Although TRADES performs comparably, PatchPro consistently delivers higher performance across various metrics. For both datasets, patch modules are applied at the network’s second-to-last layer, using a fully connected network with a single linear layer. The linear layer takes the output from the network’s penultimate layer as input, and its output is added to the original network’s output.
Answer to RQ3: PatchPro demonstrates good scalability in repairing local robustness on large-scale DNNs, which also achieves a 100% repair success rate coupled with 0% drawdown, outperforming the other tools.
TABLE V: Results of repairing local robustness on Tiny ImageNet
Model $r$ $n$ RSR/RGR/DD/DSR/DGSR/Time/s
CARE TRADE Ours CARE TRADE Ours CARE TRADE Ours CARE TRADE Ours CARE TRADE Ours CARE TRADE Ours
WRN 101-2 $\frac{2}{255}$ 500 0.0 100.0 100.0 1.2 100.0 100.0 0.1 46.8 0.0 0.0 100.0 100.0 0.0 6.6 64.4 768.7 5040.8 2029.6
1000 0.0 100.0 100.0 0.5 100.0 100.0 0.0 43.2 0.0 0.0 100.0 100.0 0.0 7.3 64.4 1345.7 7088.6 3260.7
$\frac{4}{255}$ 500 0.0 100.0 100.0 0.0 100.0 100.0 0.0 51.8 0.0 0.0 97.6 100.0 0.0 1.8 64.4 721.9 4985.5 1731.8
1000 0.0 100.0 100.0 0.0 100.0 100.0 0.1 49.5 0.0 0.0 97.2 100.0 0.0 2.1 64.4 1440.8 6976.2 3221.2
ResNet 152 $\frac{2}{255}$ 500 0.0 100.0 100.0 0.0 100.0 100.0 0.0 64.5 0.0 0.0 100.0 100.0 0.0 5.2 68.2 545.3 4667.7 2219.6
1000 0.0 100.0 100.0 0.0 100.0 100.0 0.0 62.6 0.0 0.0 99.9 100.0 0.0 6.1 68.2 1092.9 7503.5 4703.7
$\frac{4}{255}$ 500 0.0 100.0 100.0 0.0 100.0 100.0 0.0 64.5 0.0 0.0 97.6 100.0 0.0 1.3 68.2 535.4 4624.2 2336.1
1000 0.0 100.0 100.0 0.0 100.0 100.0 0.0 63.6 0.0 0.0 98.5 100.0 0.0 2.0 68.2 1028.1 7499.9 4366.8
TABLE VI: The efficacy of patch module scale
VGG19 ResNet-18
$r=4/255$ $r=8/255$ $r=4/255$ $r=8/255$
50 100 200 500 1000 50 100 200 500 1000 50 100 200 500 1000 50 100 200 500 1000
DSR/PS 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
PL 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.0 99.0 99.5 99.2 99.4 100.0 100.0 99.5 99.8 99.8
DGSR/PS 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3
PL 93.2 93.2 92.8 92.0 92.3 91.2 93.3 93.4 93.4 93.4 60.0 77.5 84.2 81.4 87.2 68.6 74.5 77.3 82.5 86.6
Time/s PS 238.7 471.6 963.0 2516.9 5441.6 229.9 470.7 950.1 2485.0 5436.0 265.8 525.1 1050.3 2586.4 5262.3 264.7 523.1 1042.8 2627.8 5325.1
PL 263.3 497.1 987.1 2562.7 4991.7 244.7 495.5 992.9 2468.7 5169.8 292.5 583.4 1155.1 3076.9 5865.6 341.5 592.4 1154 3007.7 5941.1
Verified PS
PL $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
TABLE VII: The efficacy of single patch moudle
VGG19 ResNet-18
r = 4/255 r = 8/255 r = 4/255 r = 8/255
50 100 200 500 1000 50 100 200 500 1000 50 100 200 500 1000 50 100 200 500 1000
RSR/10.0 18.0 16.5 18.2 16.7 10.0 9.0 14.5 11.8 11.6 12.0 18.0 16.5 16.2 16.2 10.0 8.0 12.0 11.6 11.3
RGR/10.0 18.0 16.5 18.2 16.7 10.0 9.0 14.5 11.8 11.6 12.0 18.0 16.5 16.2 16.2 10.0 8.0 12.0 11.6 11.3
DSR/6.0 4.0 2.5 1.8 1.3 6.0 6.0 3.0 2.0 1.3 4.0 1.0 3.5 1.8 1.3 8.0 6.0 4.0 4.8 4.1
DGSR/5.8 5.2 3.9 3.9 3.8 5.7 5.6 2.05 1.9 1.9 5.2 1.5 1.5 1.5 1.4 7.9 7.2 5.4 4.8 4.8
IV-E Impact of Patch Module Size and Quantity on Efficiency
In this experiment, we evaluate the effectiveness of patch module size and quantity in repairing local robustness properties. We perform a comparative analysis of PatchPro’s performance across different sizes and quantities, presenting the experimental results in Table VI. The performance of PatchPro is assessed using VGG19 and ResNet18 on the CIFAR-10 dataset, across various combinations of patch module scales and quantities, including small patch modules (PS) and large patch modules (PL).
Two distinct patch module sizes are considered: small and large. The small patch module consists of a single linear layer, which takes the output from the network’s penultimate layer as input. The large patch module also takes the output from the penultimate layer as input but additionally comprises three hidden layers with 200 neurons each. It is noted that larger patch networks tend to yield poorer results, possibly due to increased over approximation error when passing through ReLU nodes in the hidden layers. As the number of hidden layers increases, the cumulative over approximation error grows, leading to less effective repairs. Smaller patch scales slightly reduce the repair time. Additionally, when using DeepPoly to verify the feature layer repairs after applying smaller patches, the verification success rate is higher.
We also tested the effect of using a single patch to repair the neural network, and the results are shown in Table VII. It can be observed that the repair effectiveness using a single patch is significantly inferior to using multiple patches. A single patch only achieves 10% to 20% RSR and RGR. In contrast, using multiple patches results in both RSR and RGR reaching 100%, as shown in Table IV. Additionally, the single patch performs much worse in terms of DSR and DGSR. Therefore, using multiple patches for repair is highly beneficial.
Answer to RQ4: We observe that smaller patch modules are more effective, likely due to reduced over approximation error and shorter repair times, as well as higher verification success rates with abstract interpretation tools. For the aspect of patch module quantity, multiple patch modules significantly outperform only one patch module.
V Related work
@@
Provable DNN repair. The methods most closely related to ours are ART [25] and REASSURE [21]. In comparison to REASSURE, our approach fundamentally differs in its repair objectives. REASSURE focuses on fixing activation patterns of neural networks, yet for high-dimensional data, the input constraints of a property may encompass a large number of activation patterns. In contrast, PatchPro leverages formal verification to directly repair properties. Although ART also employs training in its repair process, it modifies parameters in the original neural network. On the other hand, PatchPro uses patch modules specifically to fix properties, ensuring the performance of the original network. Other methods such as PRDNN [20] and APRNN [22] formulate the repair problem in linear programming. However, their efficacy is limited on properties with high-dimensional polytopes, such as the robustness of image classification.
@@
Heuristic DNN repair. Utilizing heuristic algorithms such as particle swarm optimization and differential evolution, CARE [8] and Arachne [9] aim to pinpoint the neurons responsible for faults. VeRe [10] focuses on providing formal verification guidance to assist fault localization, and defines target intervals for repair synthesis. For example, DeepRepair [16] and few-shot guided mix [17] expand the set of negative samples to generate additional training data. Conversely, Tian et al. [40] augment the existing data by introducing real-world environmental effects, such as fog, to the samples. DL2 [24] fuses logical constraints and loss functions, but without convex-certified guarantees.
@@
DNN verification. Over the past decade, various approaches of DNN verification has been proposed including techniques such as constraint solving [41, 42, 43, 44], abstract interpretation [11, 45, 46, 47, 48], linear relaxation [49, 50, 51, 52, 53], global optimisation [54, 55, 56], CEGAR [57, 58, 59], reduction to two-player games [60, 61], and star-set abstraction [62, 63]. These method offer provable estimations of DNN robustness. Moreover, statistical approaches, presented in [64, 65, 66, 67, 68, 69, 70, 71, 72, 73], prove to be more efficient and scalable, particularly suited for intricate DNN structures, allowing for the establishment of quantifiable robustness at a specified confidence level. Certified training [74, 75, 76] uses convex approximation in training the loss function, but it is only used to enhance the robustness of DNN, without making it accessible to repair properties in the training procedure.
VI Conclusion
@@
We introduce PatchPro, a novel approach for property-based repair of local robustness using limited data. Our method provides patch modules as neural networks to repair within the robustness neighborhood, enabling the generalization of this defense to other inputs. In terms of efficiency, scalability, and generalization, our approach surpasses existing methods.
VII Acknowledgements
This work is supported by CAS Project for Young Scientists in Basic Research, Grant No.YSBR-040, ISCAS Basic Research ISCAS-JCZD-202302, and ISCAS New Cultivation Project ISCAS-PYFX-202201. The limitation of our work is that our patch allocation may assign inappropriate patches when the input is misclassified, and in this case PatchPro may fail to repair its local robustness. For future work, we plan to integrate verification methods based on Branch-and-Bound like [35] to enhance the performance of our approach, improving both precision and offering an alternative refinement approach. Additionally, our patch-based repair framework has the potential to evolve into a black-box repair method combined with black-box verification methods like [73].
Appendix A Proof of Theorem 3.4
Theorem 1.
Let $\varphi=(F,B(x_{i},r))$ be a local robustness property. If $\mathcal{L}(\varphi)=0$ on $B(x_{i},r)$, i.e.,
$\mathcal{L}^{*}(\varphi):=$ $\max(\mathrm{elmax}(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}+r\cdot\bm{1})+\mathrm{elmin}(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot$
$(x_{i}-r\cdot\bm{1})+\beta_{\ell},0)=0,$
@@ VII Acknowledgements
where $\mathrm{elmax}$ and $\mathrm{elmin}$ are the element-wise $\max$ and $\min$ operation, $\bm{0}$ and $\bm{1}$ are the vector in $\mathbb{R}^{n_{0}}$ with all the entries 00 and $1$, respectively, then the property $\varphi$ holds.
Proof.
@@
$\quad\quad\max_{x\in B(x_{i},r)}\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$
$=\sum_{i\in\mathbb{R}^{n_{0}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}>0\}}(\alpha_{\ell})_{i}\max_{x\in B(x_{i},r)}x$
$+\sum_{i\in\mathbb{R}^{n_{0}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}<0\}}(\alpha_{\ell})_{i}\max_{x\in B(x_{i},r)}x+\beta_{\ell}$
$=\sum_{i\in\mathbb{R}^{n_{0}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}>0\}}(\alpha_{\ell})_{i}\cdot(x_{i}+r)$
$+\sum_{i\in\mathbb{R}^{n_{0}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}<0\}}(\alpha_{\ell})_{i}\cdot(x_{i}-r)+\beta_{\ell}$
@@
$\iff$ $\sum_{\ell\neq\ell_{0}}\max(\mathrm{elmax}(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x+r\cdot\bm{1})+\mathrm{elmin}(\alpha_{\ell}^{\mathrm{T}},\bm{0})$
$\cdot(x-r\cdot\bm{1})+\beta_{\ell},0)=0$
@@
$\forall x\in B(x_{i},r),F(x)_{\ell}-F(x)_{\ell_{0}}\leq\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}\leq 0,$
@@
Appendix B An explainable example
[…]
Figure 2: A fully connected neural network $N$ with ReLU activations.
Example 1.
Consider the neural network in Fig. 2 and the inputs $x=(-0.7,1)$ labeled “2”. Within the region $B(x,0.5)$, we have its counterexample $x^{*}=(-0.2,1.5)$, which violates the local robustness. To repair the network, we need to construct a patch $P=wx$ where $w\in\mathbb{R}^{2\times 2}$.
After initializing $w$ to all $0.1$, we train the patch according the “while” loop begins at Line 7 in Alg. 1. As shown in Line 5 in Alg. 2), we first execute DeepPoly to obtain the safety violated loss function $\mathcal{L}=w_{1,1}x_{1}+w_{1,2}x_{2}-w_{2,1}x_{1}-w_{2,2}x_{2}+(0.7x_{1}+0.14x_{2}+1.08)$. To maximize $\mathcal{L}$, we set $x_{1}=-0.2$ and $x_{2}=1.5$, then we have $\mathcal{L}^{*}(w)=-0.2w_{1,1}+1.5w_{1,2}+0.2w_{2,1}-1.5w_{2,2}+1.15$. Finally, we update $w$ by $w-\eta\nabla\mathcal{L}^{*}(w)$ with the learning rate $\eta$. This process is repeated until the robustness is proven or the epochs reaches its limit $R$.
To show the subsequent repair process, we set $\eta=0.6$ and $R=1$. Under this setting, the network is not be repaired, then we need to refine the robustness to two properties by bisecting the region $B(x,0.5)$ (see Line  14–19 in Alg. 1). Specifically, by the judgment that $\partial_{1}\mathcal{L}>\partial_{2}\mathcal{L}$, we select the dimension of $x_{1}$ and divide the robustness region $B(x,0.5)$ into $[-1.2,-0.7]\times[0.5,1.5]$ and $[-0.7,-0.2]\times[0.5,1.5]$. We perform the above repair process again, and based on a more accurate abstraction provided by two new properties, we finally obtain the repaired network with the patch
$P=\begin{pmatrix}0.22&-0.8\\ -0.02&1\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}\,.$
Appendix C Limitation
As mentioned in section III-C, the adversarial examples not within any known robustness regions may lead to the establishment of incorrect new properties, and the resulting inappropriate patches allocation may ultimately affect the performance of the repaired neural network. Although this situation is rare in reality, we created an extreme setting here to investigate this weakness of our method. Specifically, considering the local robustness with radius $r=4/255$, we reuse the repaired VGG19 for CIFAR-10 and execute it over the adversarial dataset $D_{\mathrm{adv}}$ consisting of the adversarial examples generated by attacking the original network over the testset $D_{\mathrm{t}}$. The attack utilize AutoAttack with the step size of 10. By preserving the new properties established in this execution, we retest the drawdown of the repaired model over $D_{\mathrm{t}}$. The results are presented in Fig. 3, showing that the accuracy of the repaired model drops significantly under this extreme setting. This is a limitation of PatchPro, and it also echoes the statement aforementioned that adversarial detection can serve as an important supplement to our method.
Figure 3: Results under the extreme setting.
Appendix D Network architectures and accuracies of the DNN trained on MNIST